# KCL-PH-TH/2026-03

### v2

@@ Abstract
The string-inspired running vacuum model (StRVM) of inflation is based on a Chern-Simons (CS) gravity effective action, in which the only four-spacetime-derivative-order term is a gravitational anomalous CS Pontryagin density coupled to an axion. In this work, we revisit curvature-squared string-inspired effective actions, from the point of view of appropriate local field redefinitions, leaving the perturbative string scattering matrices invariant. We require simultaneously unitarity and torsion interpretation of the field strength of the Kalb-Ramond antisymmetric tensor, features characterising the (3+1)-dimensional StRVM Cosmology. Unlike the higher dimensional case, the above feature is possible in the context of (3+1)-dimensional spacetimes, obtained after string compactification. We demonstrate that the unitarity and torsion-interpretation requirements lead to a single-type of extra four-derivative terms in the effective gravitational action, not discussed in the previous literature of StRVM, which however is shown to be subleading by many orders of magnitude, compared to the terms of the StRVM framework. Hence, its presence has no practical implications for the relevant inflationary (and, hence, postinflationary) physics of the StRVM. This demonstrates the phenomenological completeness of the StRVM cosmological scenario, which is thus fully embeddable in the UV complete (quantum-gravity compatible) string theory framework.
@@ I Introduction
In a series of articles in the recent literature [1, 2, 3, 4, 5, 6, 7, 8], a model of inflation of running-vacuum-model (RVM) type [9, 10, 11, 12, 13, 14, 15, 16, 17] has been developed in the context of string theory, in which inflation is produced by a condensate of the gravitational anomaly terms of Chern-Simons (CS) type, characterizing the early Universe. This condensate is induced by primordial gravitational-wave (GW) tensor perturbations. The anomaly terms are remnants in (3+1)-spacetime dimensions of the Green-Schwarz counterterms in the string effective action, relevant to the anomaly cancellation mechanism [18]. In the string-inspired RVM model (StRVM) [2, 4] such terms are assumed present in the primordial Universe, whose dynamics is described only by fields in the massless gravitational multiplet of strings (which is also the ground state of the superstring) [19, 20]. Assuming a dynamical supergravity breaking mechanism in the very early, pre-RVM-inflation superstring Universe [2, 4, 21, 22], implies that the supersymmetry partners of the massless gravitational string multiplet become massive, with masses near the Planck scale, thereby leaving behind as relevant degrees of freedom the massless graviton, dilaton, and spin-1 antisymmetric tensor (Kalb-Ramond (KR)) fields [19], which drive the RVM inflation. In the StRVM approach the dilaton has been self consistently assumed to be stabilised to a constant value [3], which phenomenologically determines the string coupling that enters the (3+1)-dimensional gauge couplings of the low-energy effective field theory derived from the string. Thus, it plays no further rôle in the cosmology.
@@
In the StRVM cosmology so far the gravitational low-energy effective action has been assumed to be of the lowest non-trivial order [19] in an expansion in powers of the Regge slope $\alpha^{\prime}=M_{s}^{-2}$ (where $M_{s}$ is the string energy scale). The Green-Schwarz anomaly counterterms [18] enter through appropriate modifications of the field-strength three-forms $\mathcal{H}_{\mu\nu\rho}$, $\mu,\nu\,\rho=0,\dots 3$, of the spin-1 KR antisymmetric-tensor field in (3+1)-dimensions after string compactification. These modifications imply a Bianchi identity for the field-strength, which, when implemented in the path-integral of the low-energy gravitational theory, after path-integration with respect to the field strength, yields [23] an effective gravitational action term containing a massless axion-like field (KR axion) $b(x)$, coupled to the CS gravitational anomaly terms (also called Hirzebruch signature or gravitational Pontryagin density [24]). This yields a CS-gravity [25, 26] effective field theory for the primordial StRVM universe.
@@
The CS anomaly term is of quartic order in derivatives, being quadratic in curvature. When the dilaton is constant, a graviton-ghost-free, and thus unitary, Gauss-Bonnet (GB) term, which is of quadratic-order in the curvature tensors, can always be arranged to characterize the effective string action by appropriate graviton field redefinitions in perturbative string theory [27]. In a (3+1)-dimensional spacetime, the GB combination is a total derivative [24] and thus plays no rôle in the physics of the StRVM, since its spacetime integral, and hence its contribution to the effective action, vanishes, upon assuming appropriate boundary conditions.
@@
However, this may not be the end of the story. To be formally consistent with the most general effective action, expanded up to an including ${\mathcal{O}}(\alpha^{\prime})$ (four spacetime derivative) terms, one has to consider the corresponding $\mathcal{O}(\alpha^{\prime})$ terms involving the field strength of the antisymmetric tensor field, which are present in the string effective action [28, 29, 30, 31]. It is in their presence that one, then, has to perform the path integration over the $\mathcal{H}$-field strength, so as to arrive at a form of the effective action involving a canonically normalized KR axion-like field. Classically, one may naively think that such a field can arise by identifying (in (3+1) dimensions) its spacetime derivative with the Hodge dual [24] of the three form $\mathcal{H}$:
@@
with $\eta_{\mu\nu\rho\sigma}$ the Levi-Civita density (see (56), below), and replacing it in the effective action [23]. This is not quite correct though, given that we want to establish the emergence of the axion at a full quantum level, after the $\mathcal{H}$-field path integration. In principle, the fourth-order in derivatives effective action contains also quartic fields in the KR $H$-field strength, which would complicate obtaining analytic results after $\mathcal{H}$-path integration. Moreover, as we shall see in section IV, the relation (1) is modified once higher-order corrections are taken into account (cf. (90)).
@@
It is the purpose of this work to first discuss the independent effective gravitational action terms in such an expansion, after appropriate local field redefinitions, which leave the scattering (S-) matrix of perturbative string theory invariant. In this respect, we shall reexamine (using a different basis of terms) the $\mathcal{O}(\alpha^{\prime})$ terms of the bosonic part of the (Heterotic [19, 20], for concreteness and phenomenological relevance) string effective action, involving antisymmetric tensor and graviton fields [28, 29, 30, 31]. We shall impose target-space unitarity (i.e. Gauss-Bonnet-type curvature-squared terms) and torsion interpretation for the antisymmetric-torsion field strength [19, 23]. As we shall see, unlike the higher $D$-dimensional ($D>(3+1)$) case of [30, 31], in the (3+1)-dimensional effective action (after string compactification) the imposition of both requirements is compatible, which allows for a direct link with the StRVM framework [2, 4]. Then, we shall argue that there is a scheme in which the additional (compared to the StRVM action) four-derivative terms are quadratic in the $\mathcal{H}$-field, contracted appropriately with Ricci and Ricci scalar curvature tensors, and thus the $\mathcal{H}$-path integration remains Gaussian and can be performed analytically. In this way we arrive at a non-minimal coupling of $\propto\alpha^{\prime}\partial_{\mu}b\partial_{\nu}b$ terms with the Ricci tensor $R^{\mu\nu}$, as the only type of extra terms in the contorted unitary $\mathcal{O}(\alpha^{\prime})$ effective action. In cosmological settings, such terms will be shown to be extremely suppressed compared to the terms kept in the StRVM analysis [1, 2, 3, 4, 5, 6, 7, 8], and hence they do not affect the earlier conclusions on the StRVM cosmology, which is, therefore, fully embeddable in the string framework.
@@
The structure of the article is the following: in section II we formulate the the $\mathcal{O}(\alpha^{\prime})$ (four-spacetime derivatives) effective target spacetime action of the (Heterotic, for concreteness) string theory, based on the massless bosonic string multiplet of graviton and antisymmetric-tensor field strength $\mathcal{H}_{\mu\nu\rho}$, with constant dilatons. In section III we discuss field redefinitions in the ten-dimensional target spacetime of strings that leave the perturbative target spacetime scattering (S-)matrix invariant under the assumptions of unitarity and (totally antisymmetric) torsion interpretation of $\mathcal{H}_{\mu\nu\rho}$ ($\mathcal{H}$-torsion). In section IV, we verify the well-known fact of the emergence of dynamical axion-like fields in the (3+1)-dimensional effective action, after string compactification, obtained by integrating out the $\mathcal{H}$-torsion in this more general context. The so arising axions are in this sense dual to the $\mathcal{H}$-torsion in (3+1)-dimensions, as in the StRVM case, but now the (3+1)-dimensional string-inspired effective action contains extra terms, as compared to the StRVM case, In section V, we discuss the effects of these extra terms and show that they do not affect the original conclusions on the inflationary era in the StRVM. Finally, section VI contains our conclusions. Some technical aspects of our work are presented in four Appendices. Specifically, Appendix A discusses mathematical identities among generalized curvature Riemann terms, and combinations thereof, with totally antisymmetric torsion, and their decomposition into torsion-free and contorted parts, as well as other useful identities among the torsion tensors, used in our analysis. In Appendix B, we discuss, for completeness, the case where the field redefinitions are performed directly in a $D=4$ unitary effective field action, with a totally antisymmetric torsion, which may be unrelated to a string theory one, and we compare with the string-theory case. In Appendix C we make come comments on generalized contorted parity-odd curvature invariants, of potential interest to gravitational anomalies, stressing their (trivial) behaviour once the torsion is reduced to only its totally antisymmetric components, as is the case of the string-inspired $\mathcal{H}$-torsion we are examining here. Finally, Appendix D includes technical details on the contributions of the new (as compared to the study in [2, 4, 7]) terms of the ${\mathcal{O}}(\alpha^{\prime})$ effective action to the gravitational CS anomaly condensate.
@@ II The Effective Action of Heterotic String Theory
It is well known that the bosonic gravitational sector of a generic string theory consists of three massless fields [19]: a traceless, symmetric, dimensionless, spin-2 tensor field $g_{\mu\nu}$, uniquely identified with the graviton, a dimensionless spin-0 (scalar) field $\Phi$, known as the dilaton, and the dimensionless spin-1 antisymmetric (Kalb-Ramond) field $B_{\mu\nu}$. [^1] For superstrings, the above multiplet constitutes also the bosonic part of the respective ground state [19, 20]
In the closed-string sector, there is a $U(1)$ gauge symmetry $B_{\mu\nu}\rightarrow B_{\mu\nu}+\partial_{\mu}\theta_{\nu}-\partial_{\nu}\theta_{\mu}$ and, as such, the low-energy string effective action depends solely on the field strength of the Kalb-Ramond field $B_{\mu\nu}$,
@@
$\partial_{[\mu}H_{\nu\rho\sigma]}=0,$ (3)
i.e. its exterior derivative vanishes, by construction. In heterotic strings, anomaly cancellation requirements dictate that Lorentz and gauge Chern-Simons terms be added to the field strength of the Kalb-Ramond field $B_{\mu\nu}$ (the so-called Green-Schwarz (GS) mechanism [18]), such that (we use the language of differential forms [24] here, for index economy reasons):
[…]
where $\mathcal{H}_{\mu\nu\rho}=\kappa^{-1}H_{\mu\nu\rho}$ has dimensions of $mass]^{2}$, $\kappa=M_{\rm Pl}^{-1}$ is the (3+1)-dimensional gravitational coupling, with $M_{\rm Pl}=2.435\times 10^{18}~{\rm GeV}$ the reduced Planck mass, and $\alpha^{\prime}=M_{s}^{-2}$ is the Regge slope of the string, with $M_{s}$ the string mass scale, which is not necessarily equal to the Planck mass scale. The Lorentz and gauge Chern-Simons terms are
@@
respectively, where $\boldsymbol{\omega}$ is the spin connection one-form and $\boldsymbol{A}$ is the Yang-Mills gauge potential one-form. The trace Tr is over both gauge and Lorentz group indices. This modification of the Kalb-Ramond field strength leads to a corresponding modification of the Bianchi identity it satisfies, which can be written as [23]:
@@
where $\boldsymbol{F}=\boldsymbol{dA}+\boldsymbol{A}\wedge\boldsymbol{A}$ is the Yang-Mills field strength two-form and $\boldsymbol{R}^{a}{}_{b}=\boldsymbol{d\omega}^{a}{}_{b}+\boldsymbol{\omega}^{a}{}_{c}\wedge\boldsymbol{\omega}^{c}{}_{b}$ is the curvature two-form. The modified, nonzero quantity on the right-hand side of the Bianchi identity is the so-called “mixed (gauge and gravitational) quantum anomaly” [32].
We now consider the bosonic part of the heterotic effective string action up to order $\mathcal{O}(\alpha^{\prime})$. We will consider the case where the dilaton varies slowly or has been stabilized (e.g. by means of an appropriate, string-loop-induced potential) to some constant value $\Phi_{0}$, which, without loss of generality, we may set to zero. The effective string action then is:
@@
with $\nabla_{\mu}$ denoting the (torsion-free) gravitational covariant derivative, is the $\alpha^{\prime}$ perturbative correction to the action. [^4] There are $2$ more terms of the form $\nabla\mathcal{H}\nabla\mathcal{H}$ we may construct, namely $\nabla_{\gamma}\mathcal{H}_{\alpha\beta\delta}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$ and $\nabla_{\delta}\mathcal{H}_{\alpha\beta\gamma}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$. However, these can both be “converted” to $\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}$ via partial integrations modulo total derivative terms. Note that in the case of $\nabla_{\delta}\mathcal{H}_{\alpha\beta\gamma}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$ especially, we need to make use of the Bianchi identity $\nabla_{[\delta}\mathcal{H}_{\alpha\beta\gamma]}=0$ to achieve this (the reader should note at this stage that the $\mathcal{O}(\alpha^{\prime})$ corrections, due to the GS mechanism, in the Bianchi identity (7), are ignored, when the identity is implemented in the aforementioned four-derivative terms $\nabla\mathcal{H}\,\nabla\mathcal{H}$, as a result of the truncation of the respective effective action to four spacetime derivatives, which we assume throughout this work). The coefficients $\lambda_{1},\ldots\lambda_{7}$ are not uniquely identified due to field redefinition ambiguities present in the theory, which leave the perturbative S-matrix invariant [27, 28, 29, 30, 31, 23].
In what follows we shall fix these redefinitions, with the aim of arriving at the most general string effective action to order $\mathcal{O}(\alpha^{\prime})$ based on the graviton and antisymmetric tensor fields (having stabilised the dilaton). Motivated by the desire to apply it to phenomenologically realistic cosmologies, in (3+1)-dimensional spacetimes after string compactification [19], and in particular the StRVM Cosmology, we shall impose unitarity and a torsional interpretation of the totally antisymmetric tensor field strength, features which have been used in that approach [2, 3, 4, 5].
@@
In this section, we shall discuss the most general $\mathcal{O}(\alpha^{\prime})$ low-energy string effective action, compatible with the $D(=10)$-dimensional string theory (although we perform the analysis in general number of spacetime dimensions $D$, nonetheless we have in mind the low-energy theory corresponding to realistic superstring, and in particular heterotic [19], string theories). According to the equivalence theorem [33, 34, 35, 36, 37, 38], the perturbative scattering (S-)matrices of quantum field theories linked through local field redefinitions, are the same. This is what we make use of in our search for the most general effective action that describes StRVM embedded in realistic string theories, serving as consistent Ultra Violet (UV) completions of the model, compatible with quantum gravity. The equivalence theorem is valid, and thus consistent with the intuition that field redefinitions are a mere change of variables in a path integral, provided that: the redefinitions are local, invertible (i.e. non singular), and they do not affect the boundary conditions of the system, nor the spectrum of its asymptotic particle states.
The S-matrix associated with the the effective action (8) ((9), (10)) remains invariant under the following transformations of the graviton $g_{\mu\nu}$ and Kalb-Ramond $B_{\mu\nu}$ fields,

### v3

@@ Abstract
The string-inspired running vacuum model (StRVM) of inflation is based on a Chern–Simons (CS) gravity effective action in which the only four-spacetime-derivative-order term is a gravitational anomalous CS–Pontryagin density coupled to an axion. In this work, we revisit curvature-squared string-inspired effective actions from the point of view of appropriate local field redefinitions, leaving the perturbative string scattering matrices invariant. We require simultaneously unitarity and torsion interpretation of the field strength of the Kalb–Ramond antisymmetric tensor, features characterizing the (3+1)-dimensional StRVM cosmology. Unlike the higher-dimensional case, the above features are possible in the context of (3+1)-dimensional spacetimes, obtained after string compactification. We demonstrate that the unitarity and torsion interpretation requirements lead to a single type of extra four-derivative terms in the effective gravitational action, not discussed in the previous literature on StRVM, which is, however, shown to be subleading by many orders of magnitude compared to the terms of the StRVM framework. Hence, its presence has no practical implications for the relevant inflationary (and, hence, postinflationary) physics of the StRVM. This demonstrates the phenomenological completeness of the StRVM cosmological scenario, which is thus fully embeddable in the UV-complete (quantum gravity-compatible) string theory framework.
@@ I Introduction
In a series of articles in the recent literature [1, 2, 3, 4, 5, 6, 7, 8], a model of inflation of the running vacuum model (RVM) type [9, 10, 11, 12, 13, 14, 15, 16, 17] has been developed in the context of string theory, in which inflation is produced by a condensate of the gravitational anomaly terms of the Chern–Simons (CS) type, characterizing the early Universe. This condensate is induced by primordial gravitational wave (GW) tensor perturbations. The anomaly terms are remnants in (3+1)-spacetime dimensions of the Green–Schwarz counterterms in the string effective action, relevant to the anomaly cancellation mechanism [18]. In the string-inspired RVM model (StRVM) [2, 4], such terms are assumed present in the primordial Universe, whose dynamics is described only by fields in the massless gravitational multiplet of strings (which is also the ground state of the superstring) [19, 20]. Assuming a dynamical supergravity breaking mechanism in the very early, pre-RVM-inflation-superstring Universe [2, 4, 21, 22] implies that the supersymmetry partners of the massless gravitational string multiplet become massive, with masses near the Planck scale, thereby leaving behind as relevant degrees of freedom the massless graviton, dilaton, and spin-1 antisymmetric tensor (Kalb–Ramond (KR)) fields [19], which drive RVM inflation. In the StRVM approach, the dilaton is self-consistently assumed to be stabilized to a constant value [3] that phenomenologically determines the string coupling that enters the (3+1)-dimensional gauge couplings of the low-energy effective field theory derived from the string. Thus, it plays no further role in the cosmology.
@@
In StRVM cosmology, the gravitational low-energy effective action is assumed to be of the lowest non-trivial order [19] in an expansion in powers of the Regge slope $\alpha^{\prime}=M_{s}^{-2}$ (where $M_{s}$ is the string energy scale). The Green–Schwarz anomaly counterterms [18] enter through appropriate modifications of the field strength’s three forms $\mathcal{H}_{\mu\nu\rho}$, $\mu,\nu,\rho=0,\dots,3$ of the spin-1 KR antisymmetric tensor field in (3+1) dimensions after string compactification. These modifications imply a Bianchi identity for the field strength, which, when implemented in the path integral of the low-energy gravitational theory, after path-integration with respect to the field strength, yields [23], an effective gravitational action term containing a massless axion-like field (KR axion) $b(x)$ coupled to the CS gravitational anomaly terms (also called Hirzebruch signature or gravitational Pontryagin density [24]). This yields a CS-gravity [25, 26] effective field theory for the primordial StRVM Universe.
@@
The CS anomaly term is of quartic order in derivatives, being quadratic in curvature. When the dilaton is constant, a graviton-ghost-free, and thus unitary, Gauss–Bonnet (GB) term that is of quadratic-order in the curvature tensors can always be arranged to characterize the effective string action by appropriate graviton field redefinitions in perturbative string theory [27]. In a (3+1)-dimensional spacetime, the GB combination is a total derivative [24] and thus plays no role in the physics of the StRVM, since its spacetime integral, and hence its contribution to the effective action, vanishes upon assuming appropriate boundary conditions.
@@
However, this may not be the end of the story. To be formally consistent with the most general effective action, expanded up to and including ${\mathcal{O}}(\alpha^{\prime})$ (four-spacetime-derivative) terms, one has to consider the corresponding $\mathcal{O}(\alpha^{\prime})$ terms involving the field strength of the antisymmetric tensor field, which are present in the string effective action [28, 29, 30, 31]. It is in their presence that one, then, has to perform the path integration over the $\mathcal{H}$-field strength so as to arrive at a form of the effective action involving a canonically normalized KR axion-like field. Classically, one may naively think that such a field can arise by identifying (in (3+1) dimensions) its spacetime derivative with the Hodge dual [24] of the three-form $\mathcal{H}$:
@@
with $\eta_{\mu\nu\rho\sigma}$ being the Levi–Civita density (see (56), below), and replacing it in the effective action [23]. This is not quite correct though, given that we want to establish the emergence of the axion at a full quantum level after the $\mathcal{H}$-field path integration. In principle, the  fourth order in derivative effective action contains also quartic fields in the KR $H$-field strength, which would complicate obtaining analytic results after $\mathcal{H}$-path integration. Moreover, as we shall see in Section IV, the relation (1) is modified once higher-order corrections are taken into account (cf. (90)).
@@
It is the purpose of this work to first discuss the independent effective gravitational action terms in such an expansion, after  appropriate local field redefinitions, that leave the scattering (S-) matrix of perturbative string theory invariant. In this respect, we shall re-examine (using a different basis of terms) the $\mathcal{O}(\alpha^{\prime})$ terms of the bosonic part of the string effective action (heterotic [19, 20], for concreteness and phenomenological relevance), involving antisymmetric tensor and graviton fields [28, 29, 30, 31]. We shall impose target–space unitarity (i.e., Gauss–Bonnet-type curvature-squared terms) and torsion interpretation for the antisymmetric torsion field strength [19, 23]. As we shall see, unlike the higher $D$-dimensional ($D>(3+1)$) case of [30, 31], in the (3+1)-dimensional effective action (after string compactification), the imposition of both requirements is compatible, which allows for a direct link with the StRVM framework [2, 4]. Then, we shall argue that there is a scheme in which the additional (compared to the StRVM action) four-derivative terms are quadratic in the $\mathcal{H}$-field, contracted appropriately with Ricci and Ricci scalar curvature tensors, and thus the $\mathcal{H}$-path integration remains Gaussian and can be performed analytically. In this way, we arrive at a non-minimal coupling of $\propto\alpha^{\prime}\partial_{\mu}b\partial_{\nu}b$ terms, with the Ricci tensor $R^{\mu\nu}$ as the only type of extra term in the contorted unitary $\mathcal{O}(\alpha^{\prime})$ effective action. In cosmological settings, such terms will be shown to be extremely suppressed compared to the terms kept in the StRVM analysis [1, 2, 3, 4, 5, 6, 7, 8], and hence, they do not affect the earlier conclusions on StRVM cosmology, which is, therefore, fully embeddable in the string framework.
@@
The structure of this article is the following: in Section II, we formulate the $\mathcal{O}(\alpha^{\prime})$ (four-spacetime derivative) effective target spacetime action of the (heterotic, for concreteness) string theory based on the massless bosonic string multiplet of graviton and antisymmetric tensor field strength $\mathcal{H}_{\mu\nu\rho}$, with constant dilatons. In Section III, we discuss field redefinitions in the ten-dimensional target spacetime of strings that leave the perturbative target spacetime scattering (S-)matrix invariant under the assumptions of unitarity and (totally antisymmetric) torsion interpretation of $\mathcal{H}_{\mu\nu\rho}$ ($\mathcal{H}$-torsion). In Section IV, we verify the well-known fact of the emergence of dynamical axion-like fields in the (3+1)-dimensional effective action after string compactification, obtained by integrating out the $\mathcal{H}$-torsion in this more general context. The arising axions are, in this sense, dual to the $\mathcal{H}$-torsion in (3+1) dimensions, as in the StRVM case, but now, the (3+1)-dimensional string-inspired effective action contains extra terms, as compared to the StRVM case, In Section V, we discuss the effects of these extra terms and show that they do not affect the original conclusions on the inflationary era in the StRVM. Finally, Section VI contains our conclusions. Some technical aspects of our work are presented in the four appendices. Specifically, Appendix A discusses mathematical identities among generalized-curvature Riemann terms and combinations thereof, with totally antisymmetric torsion and their decomposition into torsion-free and contorted parts, as well as other useful identities among the torsion tensors used in our analysis. In Appendix B, we discuss, for completeness, the case where the field redefinitions are performed directly in a $D=4$ unitary effective field action with a totally antisymmetric torsion that may be unrelated to a string theory one, and we compare with the string theory case. In Appendix C, we make come comments on generalized contorted parity-odd curvature invariants, of potential interest to gravitational anomalies, stressing their (trivial) behavior once the torsion is reduced to only its totally antisymmetric components, as is the case of the string-inspired $\mathcal{H}$-torsion we are examining here. Finally, Appendix D includes technical details on the contributions of the new (as compared to the study in [2, 4, 7]) terms of the ${\mathcal{O}}(\alpha^{\prime})$ effective action to the gravitational CS anomaly condensate.
@@ II The Effective Action of Heterotic String Theory
It is well known that the bosonic gravitational sector of a generic string theory consists of three massless fields [19]: a traceless, symmetric, dimensionless, spin-2 tensor field $g_{\mu\nu}$, uniquely identified with the graviton; a dimensionless spin-0 (scalar) field $\Phi$, known as the dilaton; and the dimensionless spin-1 antisymmetric (Kalb–Ramond) field $B_{\mu\nu}$  [^1] For superstrings, the above multiplet constitutes also the bosonic part of the respective ground state [19, 20]..
In the closed-string sector, there is a $U(1)$ gauge symmetry $B_{\mu\nu}\rightarrow B_{\mu\nu}+\partial_{\mu}\theta_{\nu}-\partial_{\nu}\theta_{\mu}$ and, as such, the low-energy string effective action depends solely on the field strength of the Kalb–Ramond field $B_{\mu\nu}$, as follows:
@@
where the symbol $\ldots$] denotes complete antisymmetrization of the respective indices [^2] Our normalization conventions for the antisymmetrization symbol $\dots$] in indices are $\mathcal{T}_{[\mu_{1}\dots\mu_{n}]}=\frac{1}{n!}\sum_{P\in\mathcal{S}_{n}}{\rm sign}(P)\,T_{\mu_{P(1)}\mu_{P(2)}\dots\mu_{P(n)}}\,$, where the symbol $P$ denotes permutations. Odd (even) permutations have sign $(P_{\rm odd(even)})=-1(+1)$.. The Kalb–Ramond field strength, which is a three-form, satisfies the Bianchi identity as follows:
$\partial_{[\mu}H_{\nu\rho\sigma]}=0,$ (3)
In other words, its exterior derivative vanishes by construction. In heterotic strings, anomaly cancellation requirements dictate that Lorentz and gauge Chern–Simons terms be added to the field strength of the Kalb–Ramond field $B_{\mu\nu}$ (the so-called Green–Schwarz (GS) mechanism [18]), such that we use the language of differential forms [24] here, for index economy reasons:
@@
where $\mathcal{H}_{\mu\nu\rho}=\kappa^{-1}H_{\mu\nu\rho}$ has dimensions of $mass]^{2}$, $\kappa=M_{\rm Pl}^{-1}$ is the (3+1)-dimensional gravitational coupling, with $M_{\rm Pl}=2.435\times 10^{18}~{\rm GeV}$ the reduced Planck mass, and $\alpha^{\prime}=M_{s}^{-2}$ is the Regge slope of the string, with $M_{s}$ the string mass scale, which is not necessarily equal to the Planck mass scale. The Lorentz and gauge Chern–Simons terms are
@@
respectively, where $\boldsymbol{\omega}$ is the spin connection one-form and $\boldsymbol{A}$ is the Yang–Mills gauge potential one-form. The trace Tr is over both gauge and Lorentz group indices. This modification of the Kalb–Ramond field strength leads to a corresponding modification of the Bianchi identity it satisfies, which can be written as follows [23]:
@@
where $\boldsymbol{F}=\boldsymbol{dA}+\boldsymbol{A}\wedge\boldsymbol{A}$ is the Yang–Mills field strength two-form and $\boldsymbol{R}^{a}{}_{b}=\boldsymbol{d\omega}^{a}{}_{b}+\boldsymbol{\omega}^{a}{}_{c}\wedge\boldsymbol{\omega}^{c}{}_{b}$ is the curvature two-form. The modified, nonzero quantity on the right-hand side of the Bianchi identity is the so-called “mixed (gauge and gravitational) quantum anomaly” [32].
We now consider the bosonic part of the heterotic effective string action up to order $\mathcal{O}(\alpha^{\prime})$. We will consider the case where the dilaton varies slowly or has been stabilized (e.g., by means of an appropriate, string loop-induced potential) to some constant value $\Phi_{0}$, which, without loss of generality, we may set to zero. The effective string action is then as follows:
@@
where $\nabla_{\mu}$ denotes the (torsion-free) gravitational covariant derivative, and $\alpha^{\prime}$ is the perturbative correction to the action [^4] There are two more terms of the form $\nabla\mathcal{H}\nabla\mathcal{H}$ that we may construct, namely $\nabla_{\gamma}\mathcal{H}_{\alpha\beta\delta}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$ and $\nabla_{\delta}\mathcal{H}_{\alpha\beta\gamma}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$. However, these can both be “converted” to $\nabla_{\mu}\mathcal{H}^{\mu\lambda\rho}\nabla^{\nu}\mathcal{H}_{\nu\lambda\rho}$ via partial integrations of modulo total derivative terms. Note that in the case of $\nabla_{\delta}\mathcal{H}_{\alpha\beta\gamma}\nabla^{\delta}\mathcal{H}^{\alpha\beta\gamma}$ especially, we need to make use of the Bianchi identity $\nabla_{[\delta}\mathcal{H}_{\alpha\beta\gamma]}=0$ to achieve this (the reader should note at this stage that the $\mathcal{O}(\alpha^{\prime})$ corrections, due to the GS mechanism in the Bianchi identity (7), are ignored when the identity is implemented in the aforementioned four-derivative terms $\nabla\mathcal{H}\,\nabla\mathcal{H}$ as a result of the truncation of the respective effective action to four spacetime derivatives, which we assume throughout this work).. The coefficients $\lambda_{1},\ldots\lambda_{7}$ are not uniquely identified due to field redefinition ambiguities present in the theory, which leave the perturbative S-matrix invariant [27, 28, 29, 30, 31, 23].
In what follows, we shall fix these redefinitions with the aim of arriving at the most general string effective action to order $\mathcal{O}(\alpha^{\prime})$ based on the graviton and antisymmetric tensor fields (having stabilized the dilaton). Motivated by the desire to apply it to phenomenologically realistic cosmologies, in (3+1)-dimensional spacetimes after string compactification [19], and in particular the StRVM cosmology, we shall impose unitarity and a torsional interpretation of the totally antisymmetric tensor field strength, features that have been used in that approach [2, 3, 4, 5].
@@
In this section, we shall discuss the most general $\mathcal{O}(\alpha^{\prime})$ low-energy string effective action, compatible with the $D(=10)$-dimensional string theory (although we perform the analysis in a general number of spacetime dimensions $D$, we have in mind nonetheless the low-energy theory corresponding to realistic superstring, and in particular heterotic [19], string theories). According to the equivalence theorem [33, 34, 35, 36, 37, 38], the perturbative scattering (S-)matrices of quantum field theories linked through local field redefinitions are the same. This is what we make use of in our search for the most general effective action that describes StRVM embedded in realistic string theories, serving as consistent Ultra Violet (UV) completions of the model, compatible with quantum gravity. The equivalence theorem is valid, and thus consistent with the intuition that field redefinitions are a mere change in variables in a path integral, provided that the redefinitions are local, invertible (i.e., non singular), and do not affect the boundary conditions of the system, nor the spectrum of its asymptotic particle states.
The S-matrix associated with the the effective action (8) ((9), (10)) remains invariant under the following transformations of the graviton $g_{\mu\nu}$ and Kalb–Ramond $B_{\mu\nu}$ fields: